OpenStudy (anonymous):

Proof: (cosx-sinx)^2=1-sin2x

geerky42 (geerky42):

\[(\cos x-\sin x)^2 = \cos^2 x - 2\cos x\sin x + \sin^2 x\]Just remember that \(\cos^2x+\sin^2x=1\) So you have \(1- 2\cos x\sin x\) \(2\cos x\sin x\) reminds me of something, like Double Angle Formulas?

OpenStudy (anonymous):

http://www.sosmath.com/trig/Trig5/trig5/trig5.html

OpenStudy (anonymous):

I know that sin2x=2sinxcosx

OpenStudy (anonymous):

well so now now u have (1- sin2x) so ur proof is done

OpenStudy (anonymous):

Ok thanks!